The Tomas School of Falconry is considering replacing a piece of equipment used in making falconry perches. The existing machine has 3 years useful life remaining. The machine had an original cost of $50,000, 3 years ago, and is projected to have a salvage value of $5,000 in 3 years time. The existing machine has a current market value of $30,000. The company has been depreciating the machine down to its salvage value using straight line depreciation. The new machine has a useful life of 3 years, will cost $75,000, and will be depreciated using the MACRS 3 year table (yr-1 33.33%, yr-2 44.44%, yr-3 14.82%, yr-4 7.41%). The new machine will have an expected salvage value of $8,000 in 3 years time. The new machine will save $40,000 a year in costs. There will be no change in Net Working Capital usage for the new machine versus the old machine. The company is in the 40% tax bracket. The Tomas School of Falconry has 12 million shares of stock outstanding with a current market price of $11.30. The stock’s beta is 2.3. The return on the S&P500 is expected to be 11%, while the return on 1 year Treasury bills is 2.2%. The company has $30 million in face value of debt outstanding. The debt has a price of 93’21. The bonds have a coupon rate of 7%, and pay coupons semi-annually. The bonds will mature in 25 years. Based on NPV should the School replace the existing machine?
Answer #1
First we will find the weighted average cost of capital to discount the cash flows :
Cost of equity
Given the details, we will find the cost of equity as per CAPM,
ie.ke=RFR+(Beta*Market risk premium)
where Market Risk Premium=Stock's Market return-Risk Free Rate
so, ke=2.2%+(2.3*(11%-2.2%))=
22.44%
Cost of bonds:
Using the formula to find the present value of bonds,
PV/Price of the bond=(Pmt.*(1-(1+r)^-n)/r)+(FV/(1+r)^n)
where price of the bond is given to be $ 93.21 per $ 100 FV
Pmt.= the semi-annual coupon pmt.= 7%/2*100= $ 3.5 per $ 100 FV
r= semi-annual yield /cost which we need to find--- ??
n= no.of semi-annual coupon periods= 25*2= 50
FV = Face value , which we are taking as $ 100
So, plugging in the values in the formula,
93.21=(3.5*(1-(1+r)^-50)/r)+(100/(1+r)^50)
& solving for r, we get the semi-annual before-tax yield/cost as
3.81%
Now the annual before-tax cost=(1+3.81%)^2-1=
7.77%
so, the annual after-tax cost=BT cost*(1-Tax Rate)
ie.7.77%*(1-40%)=
4.66%
Now, we have the cost of equity as 22.44% &
the cost of debt/bonds as 4.66%
we need to calculate the market value weights of equity & debt
Capital Mkt.values Wt.to Total Cost Wt.*Cost
Equity 12000000*11.30= 135600000 82.90% 22.44% 18.60%
Debt 30000000*93.21/100= 27963000 17.10% 4.66% 0.80%
Total 163563000 100.00% 19.40%
SO, WACC to discount the cash flows = 19.40%
NPV analysis of the Replacement decision:
Cost of new m/c -75000
After-tax sale value of existing m/c(ref. wkgs.1) 29000
PV of after-tax savings in opg. Costs(40000*(1-40%)*2.12643)-----P/A,i=19.40%,n=3 yrs. 51034
PV of Incl. dep.tax shields(Ref. wkgs. 2) 12865
PV of after-tax sale(at end yr.6) of old m/c lost (-5000*(1-40%)/1.1940^3) -1762
PV of after-tax salvage of new m/c(ref. wkgs.3)7023/1.194^3= 4126
NPV of the replacement decision 20263
As the NPV of the replacement decision is POSITIVE,it is RECOMMENDED to be replaced.
Workings:
1.After-tax Sale value of existing m/c
Cost 3 years ago 50000
Acc. Depn.(50000-5000)/6*3 22500
Carrying value,now 27500
Sale value,now 30000
Gain on sale 2500
Tax on gain 1000
so,After-tax sale cash flow for existing m/c 29000
2.Annual depn.
Year Existing m/c New m/c Incl.depn. Incl.Tax shield on depn. At 40% PV of incl. dep.tax sh. At 19.40%
1 7500 75000*33.33%= 24998 17498 6999 5861.81
2 7500 75000*44.44%= 33330 25830 10332 7247.29
3 7500 75000*14.82%%= 11115 3615 1446 849.48
4 75000*7.41% -5558 -5558 -2223 -1093.76 ####
12864.82
#### Depn. Tax shield for yr. 4 lost due to sale at end of Yr. 3
3.After-tax Sale value of new m/c
Cost 75000
Acc. Depn.(75000*(33.33+44.44+14.82)%) 69443
Carrying value 5558
Sale value 8000
Gain on sale 2443
Tax on gain 977
so,After-tax sale cash flow for new m/c 7023